# Properties of kaon-condensed phase in hyperon-mixed matter with three-baryon forces

### v1

@@ Abstract
Coexistent phase of kaon condensates and hyperons [($Y$+$K$) phase] in beta equilibrium with electrons and muons is investigated as a possible form of dense hadronic phase with multi-strangeness. The effective chiral Lagrangian for kaon-baryon and kaon-kaon interactions is utilized within chiral symmetry approach in combination with the interaction model between baryons. For the baryon-baryon interactions, we adopt the minimal relativistic mean-field theory with exchange of scalar mesons ($\sigma$, $\sigma^{\ast}$) and vector mesons ($\omega$, $\rho$, $\phi$) between baryons without including the nonlinear self-interacting meson field terms. In addition, the universal three-baryon repulsion and the phenomenological three-nucleon attraction are introduced as density-dependent effective two-body potentials.
@@ I Introduction
Toward understanding a realistic form of hadronic matter, we investigated a possible coexistence of KC and hyperons $Y$ ($Y$=$\Lambda$, $\Sigma^{-}$, $\Xi^{-}$, $\cdots$) in the ground state of neutron-star matter [abbreviated to ($Y$+$K$) phase ] m2008 ; mmtt2019 ; mmt2021 . In the ($Y$+$K$) phase, softening of the EOS results from combined effects of decreasing energy by the $s$-wave $K$-baryon ($B$) attraction and avoiding the $N$-$N$ repulsion at high densities by hyperon admixture nyt02 . Although the onset density of KC is in general pushed up to a higher density in the hyperon-mixed matter than the case of nucleon matter ekp95 ; kpe95 ; sm96 , the energy gain becomes large once KC occurs in the hyperon-mixed matter. In Ref. m2008 , both kaon condensates and hyperon-mixing give large energy gain, leading to a local minimum of the energy with respect to $\rho_{\rm B}$, which suggests self-bound objects with ($Y+K$) phase. However, such EOS including the local energy minimum cannot reproduce massive compact stars as large as 2 $M_{\odot}$, so that the self-bound star with ($Y$+$K$) phase is unlikely to exist. Most of the other works on coexistent phase of KC and hyperon-mixed matter were not able to exceed the maximum mass as large as 1.85 $M_{\odot}$ m2008 except for the case of density-dependent relativistic mean-field (RMF) theory, where specific density-dependence of the coupling constants between mediating mesons and baryons are assumed cb2014 ; ts2020 ; mbb2021 .
@@
Recently, theoretical approaches in the chiral effective field theory ($\chi$EFT) have been applied for the $YNN$ interaction petschauer2020 ; kohno2018 . As for $\Lambda$ in matter, a possibility of the strongly repulsive $\Lambda$ potential at high density and resulting suppression of the $\Lambda$-mixing has been suggested with reference to the $\chi$EFT kohno2018 ; gkw2020 ; jinno2023 . It has also been pointed out that the $\Lambda$-suppression at high densities is consistent with the $\Lambda$ directed flow data in heavy-ion collisions nara2022 . Therefore there is still controversy about the $\Lambda$ potential at high densities. It is expected that the problem is to be resolved by forthcoming hypernuclear experiments taskforce2022 . In this paper, we base our $B$-$B$ interaction model on the RMF theory, where $\Lambda$ hyperons necessarily appear at $\rho_{\rm B}$ = (2.5--3.0) $\rho_{0}$ with the nuclear saturation density $\rho_{0}$ (= 0.16 fm-3).
@@
In the early stages of a series of our works, we considered the ($Y$+$K$) phase by the use of the interaction model where $K$-$B$ and $K$-$K$ interactions are described by the effective chiral Lagrangian (abbreviated to ChL) and the $B$-$B$ interactions are given by exchange of mesons [$\sigma$, $\sigma^{\ast}$ ($\sim\bar{s}s$) for scalar mesons, $\omega$, $\rho$, $\phi$ ($\sim\bar{s}\gamma^{0}s$ on the assumption of an ideal mixing between $\omega$ and $\phi$ mesons) for vector mesons] in the RMF with the nonlinear self-interacting (NLSI) meson fields mmtt2019 ; mmt2022 ; bb77 . The NLSI terms generate many-baryon forces through the equations of motion for the meson mean fields. However, they are introduced solely to reproduce the ground state properties such as the incompressibility and slope of the symmetry energy around the saturation density. It has been shown that the NLSI terms are not relevant to repulsive energy corresponding to the UTBR leading to the stiff EOS at high densities and that it cannot be compensated with large attractive energy due to the appearance of the ($Y$+$K$) phase mmt2022 .
@@
In order to keep from the significant softening of the EOS with the ($Y$+$K$) phase, we have proposed another prescription for introducing the UTBR. We adopt the RMF model for two-body $B$-$B$ interaction mediated by meson-exchange, removing the NLSI meson potentials. We call this model a “minimal RMF” (abbreviated to MRMF) throughout this paper. Instead of the NLSI terms, we introduce the density-dependent effective two-body potentials for the UTBR, which has been derived from the string-junction model by Tamagaki t2008 (SJM2) and originally applied to $Y$-mixed matter by Tamagaki, Takatsuka and Nishizakitnt2008 . Together with the UTBR, phenomenological three-nucleon attraction (TNA) has been taken into account, and we have obtained the baryon interaction model that reproduces saturation properties of symmetric nuclear matter (SNM) together with empirical values of incompressibility and symmetry energy at $\rho_{0}$. In addition, by specifying the slope $L$ of the symmetry energy within an acceptable range, we set stiffness of the EOS coming from two-baryon repulsion (Two-BR) through the vector meson exchange. Then we have investigated the ($Y$+$K$) phase based upon the ChL coupled with this baryon interaction model (MRMF+UTBR+TNA) mmt2021 . It has been shown that the EOS and the resulting gravitational mass and radius of compact stars with the ($Y$+$K$) phase are consistent with recent observations of massive neutron stars.
@@
In this paper, we further elucidate the onset mechanisms of KC in hyperon-mixed matter and the EOS with the ($Y$+$K$) phase in detail with the use of the (ChL+MRMF+UTBR+TNA) model, following up the results of Ref. mmt2021 . Primary points of view being addressed in the paper are interplay between KC and hyperons through chiral dynamics in dense hadronic matter and combined effect of the UTBR and Two-BR on the stiffness of the EOS at high densities.
@@
The paper is organized as follows. In Sec. II, chiral symmetry approach for kaon condensation based on the effective chiral Lagrangian is introduced. The “$K$-$B$ sigma terms” are estimated by taking into account of the nonlinear effect with respect to the strange quark mass beyond chiral perturbation in the next to leading order. In Sec. III, the MRMF is explained in meson-exchange picture for $B$-$B$ interaction. In addition, the UTBR (SJM2 as a specific model) and TNA are introduced phenomenologically. Section IV gives description of the ground state for the ($Y$+$K$) phase: energy density expression for the ($Y$+$K$) phase, together with classical field equations for KC and meson mean fields, and ground state conditions. In Sec. V, the results on the properties of symmetric nuclear matter and pure neutron matter with our interaction model are given, prior to the results on KC in hyperon-mixed matter. The dependence of the EOS on the slope $L$ is also discussed. Section VI is devoted to the description of pure hyperon-mixed matter, which is followed by the discussion of kaon properties in hyperon-mixed matter and onset of KC. In Sec. VII, results of the ground state properties of the ($Y$+$K$) phase and the EOS are given. In Sec. VIII, static properties of neutron stars with the ($Y$+$K$) phase are discussed with the results such as gravitational mass to radius, mass to baryon number density relations, and density distribution in the core with the ($Y$+$K$) phase for typical masses of neutron stars. Summary and concluding remarks are given in Sec. IX. In Appendix X A, the expressions of the baryon masses and the $K$-$B$ sigma terms are given by the use of the Feynman-Hellmann theorem for the effective chiral Lagrangian. In Appendix X B, the expressions for the $s$-wave on-shell $K$-$N$ scattering lengths are given in order to determine the coefficients of the range terms and coupling constants associated with the $\Lambda$ (1405). The expressions for quantities related to saturation properties in SNM, i. e. the incompressibility, the symmetry energy, and the slope $L$ of the symmetry energy, are given in Appendix X C.
@@ II Chiral symmetry approach for kaon condensation
The ($Y$+$K$) phase is composed of kaon condensates and hyperon-mixed baryonic matter together with leptons, being kept in beta equilibrium, charge neutrality, and baryon number conservation. In the following, we simply take into account protons, neutrons, $\Lambda$, $\Sigma^{-}$, and $\Xi^{-}$ hyperons for baryons and electrons and muons for leptons.
@@ II.1 Kaon-baryon and multi-kaon interactions
We base our model for $K$-$B$ and $K$-$K$ interactions upon the effective chiral SU(3)L×$\times$×SU(3)R Lagrangian, which is given by
@@
where the first and second terms on the r. h. s. of Eq. (11) are the kinetic and mass terms of the kaons $U$ in the nonlinear representation, respectively, and $\Lambda_{\chi{\rm SB}}$ is the chiral symmetry breaking scale ($\sim$ 1 GeV), $M$ the quark mass matrix, $M\equiv{\rm diag}(m_{u},m_{d},m_{s})$ with the quark masses $m_{i}$kn86 . The nonlinear charged kaon field $U$ is represented by a chiral rotation of the meson vacuum ${\bf 1}_{\rm vac}$ as $U=g_{\rm L}{\bf 1}_{\rm vac}g_{\rm R}^{\dagger}=\xi^{2}$ with $g_{\rm L}=g_{\rm R}^{\dagger}=\xi=\exp[\sqrt{2}i(K^{+}T_{4+i5}+K^{-}T_{4-i5})/f]$, where $T_{4\pm i5}$ ($\equiv T_{4}\pm iT_{5}$) is the flavor SU(3) generators and $f$ is the meson decay constant. The value of $f$ is set to be the one of the pion decay constant ($f_{\pi}$ = 93 MeV) in the lowest-order in chiral perturbation theory. The kaon rest mass is identified with $m_{K}=[\Lambda_{\chi{\rm SB}}(m_{u}+m_{s})]^{1/2}$ and is set to be the empirical value (=493.677 MeV).
The third term in Eq. (11) denotes the kinetic and mass terms for the octet baryon field $\Psi$, where $M_{\rm B}$ is a baryon mass generated by spontaneous chiral symmetry breaking in the chiral limit. The fourth term in Eq. (11) gives the $s$-wave $K$-$B$ vector interaction corresponding to the Tomozawa-Weinberg term with $V_{\mu}$ being the vector current of the kaons defined by $V_{\mu}\equiv i/2(\xi^{\dagger}\partial_{\mu}\xi+\xi\partial_{\mu}\xi^{\dagger})$. The fifth and sixth terms are the $K$-$B$ axial-vector interaction with the mesonic axial-vector current defined by $A_{\mu}\equiv i/2(\xi^{\dagger}\partial_{\mu}\xi-\xi\partial_{\mu}\xi^{\dagger})$. Throughout this paper, we simply omit these axial-vector coupling terms and retain only the $s$-wave $K$-$B$ vector interaction in order to figure out the consequences from $s$-wave kaon condensation.
The remaining terms in Eq. (11) give the terms in the next to leading-order $O(m_{K}^{2})$ in chiral expansion. The terms with coefficients $a_{1}\sim a_{3}$ bring about the $s$-wave $K$-$B$ scalar interaction which explicitly breaks chiral symmetry. The last two terms with $d_{1}$ and $d_{2}$ are range terms. In addition to (11), we further take into account phenomenologically the pole term of the $\Lambda$ (1405) (abbreviated to $\Lambda^{\ast}$), which lies by $\sim$ 30 MeV below the $K^{-}p$ threshold. The $\Lambda^{\ast}$ pole contribution is introduced by the Lagrangian for $K^{-}p\Lambda^{\ast}$ interaction:
@@
In Eq. (18), the dimensionless constants $d_{b}$ ($b=p,n,\Lambda,\Sigma^{-},\Xi^{-}$) of the range terms are related with $d_{p}\equiv(d_{1}+d_{2})m_{K}/2$ = $d_{\Xi^{-}}$, $d_{n}\equiv d_{1}m_{K}/2$ = $d_{\Sigma^{-}}$, $d_{\Lambda}=(d_{1}+5d_{2}/6)m_{K}/2$ in terms of $d_{1}$, $d_{2}$, and $m_{K}$. The mass difference between the $\Lambda^{\ast}$ and $p$ is denoted as $\delta M_{\Lambda^{\ast}p}\equiv M_{\Lambda^{\ast}}-M_{p}$, and $\gamma_{\Lambda^{\ast}}$ is introduced as the imaginary part of the inverse propagator for the $\Lambda^{\ast}$. The constants $d_{p}$, $d_{n}$, $g_{\Lambda^{\ast}}$, and $\gamma_{\Lambda^{\ast}}$ are determined to reproduce the $s$-wave on-shell $K$-$N$ scattering lengths including the pole contribution from $\Lambda^{\ast}$ martin1981 . 111 There has been a novel improvement on the $\bar{K}$-$N$ interaction by the SIDDHARTA experiment through the kaonic hydrogen $X$-rays siddharta2011 , while there is not a quantitatively large difference of the value of the $s$-wave $\bar{K}N$ scattering length between Ref. martin1981 and siddharta2011 . Therefore we adopt the conventional values in Ref. martin1981 . One obtains $g_{\Lambda^{\ast}}$ = 0.583, $\gamma_{\Lambda^{\ast}}$=12.4 MeV, $d_{p}$ = $0.351-\Sigma_{Kp}/m_{K}$, $d_{n}$ = $0.130-\Sigma_{Kn}/m_{K}$ (see Appendix X.2). As is the case with Ref.  fmmt1996 , the range terms and $\Lambda^{\ast}$-pole term are absorbed into the effective baryon mass as shown in Eq. (18).
It has become established by the use of chiral SU(3) dynamics that the $\Lambda^{\ast}$ is the $\bar{K}N$ bound state with the strong couplings such as $\bar{K}N\leftrightarrows\pi\Sigma$ ksw1995 ; or1998 ; lk2002 ; ihw2011 ; ro2000 . Properties of $\bar{K}$$N$ dynamics in the medium was studied in the chiral unitary approach by taking into account the self-consistent $\bar{K}N$ self-energy and the Pauli-blocking effects for the nucleon in the medium ro2000 . The two resonance pole structure of the $\Lambda^{\ast}$ has also been advocated om2001 ; joo2003 ; hj2012 . In this paper, we simply take a conventional single pole picture for the $\Lambda^{\ast}$, since, as shown in Sec. VI.3, the $\Lambda^{\ast}$ contribution to the off-mass shell $K^{-}$ self-energy for $\rho_{\rm B}\gtrsim\rho_{0}$ is negligible, irrespective of details of the $\Lambda^{\ast}$ structure.

### v2

# Properties of a kaon-condensed phase in hyperon-mixed matter with three-baryon forces

@@ Abstract
Coexistent phase of kaon condensates and hyperons [($Y$+$K$) phase] in beta equilibrium with electrons and muons is investigated as a possible form of dense hadronic phase with multi-strangeness. The effective chiral Lagrangian for kaon-baryon and kaon-kaon interactions is utilized within chiral symmetry approach in combination with the interaction model between baryons. For the baryon-baryon interactions, one adopts the minimal relativistic mean-field theory with exchange of scalar mesons ($\sigma$, $\sigma^{\ast}$) and vector mesons ($\omega$, $\rho$, $\phi$) between baryons without including the nonlinear self-interacting meson field terms. In addition, the universal three-baryon repulsion and the phenomenological three-nucleon attraction are introduced as density-dependent effective two-body potentials.
@@ I Introduction
Toward understanding a realistic form of hadronic matter, a possible coexistence of KC and hyperons $Y$ ($Y$=$\Lambda$, $\Sigma^{-}$, $\Xi^{-}$, $\cdots$) in the ground state of neutron-star matter [abbreviated to ($Y$+$K$) phase ] has been investigated m2008 ; mmtt2019 ; mmt2021 . In the ($Y$+$K$) phase, softening of the EOS results from combined effects of decreasing energy by the $s$-wave $K$-baryon ($B$) attraction and avoiding the $N$-$N$ repulsion at high densities by hyperon admixture nyt02 . Although the onset density of KC is in general pushed up to a higher density in the hyperon-mixed matter than the case of nucleon matter ekp95 ; kpe95 ; sm96 , the energy gain becomes large once KC occurs in the hyperon-mixed matter. In Ref. m2008 , both kaon condensates and hyperon-mixing give large energy gain, leading to a local minimum of the energy with respect to $\rho_{\rm B}$, which suggests self-bound objects with ($Y+K$) phase. However, such EOS including the local energy minimum cannot reproduce massive compact stars as large as 2 $M_{\odot}$, so that the self-bound star with ($Y$+$K$) phase is unlikely to exist. Most of the other works on coexistent phase of KC and hyperon-mixed matter were not able to exceed the maximum mass as large as 1.85 $M_{\odot}$ m2008 except for the case of density-dependent relativistic mean-field (RMF) theory, where specific density-dependence of the coupling constants between mediating mesons and baryons are assumed cb2014 ; ts2020 ; mbb2021 .
@@
Recently, theoretical approaches in the chiral effective field theory ($\chi$EFT) have been applied for the $YNN$ interaction petschauer2020 ; kohno2018 . As for $\Lambda$ in matter, a possibility of the strongly repulsive $\Lambda$ potential at high density and resulting suppression of the $\Lambda$-mixing has been suggested with reference to the $\chi$EFT kohno2018 ; gkw2020 ; jinno2023 . It has also been pointed out that the $\Lambda$-suppression at high densities is consistent with the $\Lambda$ directed flow data in heavy-ion collisions nara2022 . Therefore there is still controversy about the $\Lambda$ potential at high densities. It is expected that the problem is to be resolved by forthcoming hypernuclear experiments taskforce2022 . In this paper, the $B$-$B$ interaction model is based on the RMF theory, where $\Lambda$ hyperons necessarily appear at $\rho_{\rm B}$ = (2.5--3.0) $\rho_{0}$ with the nuclear saturation density $\rho_{0}$ (= 0.16 fm-3).
@@
In the early stages of a series of our works, we considered the ($Y$+$K$) phase by the use of the interaction model where $K$-$B$ and $K$-$K$ interactions are described by the effective chiral Lagrangian (abbreviated to ChL) and the $B$-$B$ interactions are given by exchange of mesons [$\sigma$, $\sigma^{\ast}$ ($\sim\bar{s}s$) for scalar mesons, $\omega$, $\rho$, $\phi$ ($\sim\bar{s}\gamma^{0}s$ on the assumption of an ideal mixing between $\omega$ and $\phi$ mesons) for vector mesons] in the RMF with the nonlinear self-interacting (NLSI) meson fields mmtt2019 ; mmt2022 ; bb77 . As is discussed in Ref. sw1997 , many-body effects, in particular, all of the leading terms at each power of baryon density arising from the self-consistency condition are repulsive. The NLSI terms generate many-baryon forces through the equations of motion for the meson mean fields. However, they are introduced solely to reproduce the ground state properties of symmetric nuclear matter (SNM) such as the incompressibility and slope of the symmetry energy around the saturation density. With regard to saturation mechanisms of the SNM, it has been pointed out that large cancellation between the repulsive NLSI terms and two-body attraction maintains the saturation energy at $\rho_{0}$ mmt2022 . [See also Appendix X.4.] Therefore, the many-baryon forces induced by the NLSI terms lead to quite different picture for the saturation mechanisms compared to the case of the conventional nuclear matter theory, where two-body attraction has a main contribution to the saturation energy and the three-nucleon attraction (TNA) and three-nucleon repulsion (TNR) have complementary contribution to the energy. But both TNA and TNR play an indispensable role below and beyond $\rho_{0}$, respectively, in order to reproduce the empirical saturation point of the SNM lp1981 ; wiringa1988 ; apr1998 . Furthermore, it has been shown that the NLSI terms are not relevant to repulsive energy corresponding to the UTBR leading to the stiff EOS at high densities and that it cannot be compensated with large attractive energy due to the appearance of the ($Y$+$K$) phase mmt2022 .
@@
In order to keep from the significant softening of the EOS with the ($Y$+$K$) phase, another prescription for introducing many-baryon forces has been proposed, standing on the conventional nuclear matter theory around the saturation density. One adopts the RMF model for two-body $B$-$B$ interaction mediated by meson-exchange, removing the NLSI meson potentials. One calls this model a “minimal RMF” (abbreviated to MRMF) throughout this paper. Instead of the NLSI terms, one introduces the density-dependent effective two-body potentials for the UTBR, which has been derived from the string-junction model by Tamagaki t2008 (SJM2) and originally applied to $Y$-mixed matter by Tamagaki, Takatsuka and Nishizakitnt2008 . Together with the UTBR, phenomenological three-nucleon attraction (TNA) has been taken into account, and one has obtained the baryon interaction model that reproduces saturation properties of SNM together with empirical values of incompressibility and symmetry energy at $\rho_{0}$. In addition, by specifying the slope $L$ of the symmetry energy within an acceptable range, one sets stiffness of the EOS coming from two-baryon repulsion (Two-BR) through the vector meson exchange. Then one has investigated the ($Y$+$K$) phase based upon the ChL coupled with this baryon interaction model (MRMF+UTBR+TNA) mmt2021 . It has been shown that the EOS and the resulting gravitational mass and radius of compact stars with the ($Y$+$K$) phase are consistent with recent observations of massive neutron stars.
@@
In this paper, one further elucidates the onset mechanisms of KC in hyperon-mixed matter and the EOS with the ($Y$+$K$) phase in detail with the use of the (ChL+MRMF+UTBR+TNA) model, following up the results of Ref. mmt2021 . Primary points of view being addressed in the paper are interplay between KC and hyperons through chiral dynamics in dense hadronic matter and combined effect of the UTBR and Two-BR on the stiffness of the EOS at high densities.
@@
The paper is organized as follows. In Sec. II, chiral symmetry approach for kaon condensation based on the effective chiral Lagrangian is introduced. The “$K$-$B$ sigma terms” are estimated by taking into account of the nonlinear effect with respect to the strange quark mass beyond chiral perturbation in the next to leading order. In Sec. III, the MRMF is explained in meson-exchange picture for $B$-$B$ interaction. In addition, the UTBR (SJM2 as a specific model) and TNA are introduced phenomenologically. Section IV gives description of the ground state for the ($Y$+$K$) phase: energy density expression for the ($Y$+$K$) phase, together with classical field equations for KC and meson mean fields, and ground state conditions. In Sec. V, the results on the properties of symmetric nuclear matter and pure neutron matter with the present interaction model are given, prior to the results on KC in hyperon-mixed matter. The dependence of the EOS on the slope $L$ is also discussed. Section VI is devoted to the description of pure hyperon-mixed matter, which is followed by the discussion of kaon properties in hyperon-mixed matter and onset of KC. In Sec. VII, results of the ground state properties of the ($Y$+$K$) phase and the EOS are given. In Sec. VIII, static properties of neutron stars with the ($Y$+$K$) phase are discussed with the results such as gravitational mass to radius, mass to baryon number density relations, and density distribution in the core with the ($Y$+$K$) phase for typical masses of neutron stars. Summary and concluding remarks are given in Sec. IX. In Appendix X.1, the expressions of the baryon masses and the $K$-$B$ sigma terms are given by the use of the Feynman-Hellmann theorem for the effective chiral Lagrangian. In Appendix X.2, the expressions for the $s$-wave on-shell $K$-$N$ scattering lengths are given in order to determine the coefficients of the range terms and coupling constants associated with the $\Lambda$ (1405). The expressions for quantities related to saturation properties in SNM, i. e. the incompressibility, the symmetry energy, and the slope $L$ of the symmetry energy, are given in Appendix X.3. In Appendix X.4, the RMF model with the NLSI terms [(MRMF+NLSI) model] is overviewed for comparison with the (MRMF+UTBR+TNA) model.
@@ II Chiral symmetry approach for kaon condensation
The ($Y$+$K$) phase is composed of kaon condensates and hyperon-mixed baryonic matter together with leptons, being kept in beta equilibrium, charge neutrality, and baryon number conservation. In the following, one simply takes into account protons, neutrons, $\Lambda$, $\Sigma^{-}$, and $\Xi^{-}$ hyperons for baryons and electrons and muons for leptons.
@@ II.1 Kaon-baryon and multi-kaon interactions
The model for $K$-$B$ and $K$-$K$ interactions is based upon the effective chiral SU(3)L×$\times$×SU(3)R Lagrangian, which is given by
@@
where the first and second terms on the r. h. s. of Eq. (11) are the kinetic and mass terms of the kaons $U$ in the nonlinear representation, respectively, and $\Lambda_{\chi{\rm SB}}$ is the chiral symmetry breaking scale ($\approx$ 1 GeV), $M$ the quark mass matrix, $M\equiv{\rm diag}(m_{u},m_{d},m_{s})$ with the quark masses $m_{i}$kn86 . The nonlinear charged kaon field $U$ is represented by a chiral rotation of the meson vacuum ${\bf 1}_{\rm vac}$ as $U=g_{\rm L}{\bf 1}_{\rm vac}g_{\rm R}^{\dagger}=\xi^{2}$ with $g_{\rm L}=g_{\rm R}^{\dagger}=\xi=\exp[\sqrt{2}i(K^{+}T_{4+i5}+K^{-}T_{4-i5})/f]$, where $T_{4\pm i5}$ ($\equiv T_{4}\pm iT_{5}$) is the flavor SU(3) generators and $f$ is the meson decay constant. The value of $f$ is set to be the one of the pion decay constant ($f_{\pi}$ = 93 MeV) in the lowest-order in chiral perturbation theory. The kaon rest mass is identified with $m_{K}=[\Lambda_{\chi{\rm SB}}(m_{u}+m_{s})]^{1/2}$ and is set to be the empirical value (=493.677 MeV).
The third term in Eq. (11) denotes the kinetic and mass terms for the octet baryon field $\Psi$, where $M_{\rm B}$ is a baryon mass generated by spontaneous chiral symmetry breaking in the chiral limit. The fourth term in Eq. (11) gives the $s$-wave $K$-$B$ vector interaction corresponding to the Tomozawa-Weinberg term with $V_{\mu}$ being the vector current of the kaons defined by $V_{\mu}\equiv i/2(\xi^{\dagger}\partial_{\mu}\xi+\xi\partial_{\mu}\xi^{\dagger})$. The fifth and sixth terms are the $K$-$B$ axial-vector interaction with the mesonic axial-vector current defined by $A_{\mu}\equiv i/2(\xi^{\dagger}\partial_{\mu}\xi-\xi\partial_{\mu}\xi^{\dagger})$. Throughout this paper, one simply omits these axial-vector coupling terms and retain only the $s$-wave $K$-$B$ vector interaction in order to figure out the consequences from $s$-wave kaon condensation.
The remaining terms in Eq. (11) give the terms in the next to leading-order $O(m_{K}^{2})$ in chiral expansion. The terms with coefficients $a_{1}\sim a_{3}$ bring about the $s$-wave $K$-$B$ scalar interaction which explicitly breaks chiral symmetry. The last two terms with $d_{1}$ and $d_{2}$ are range terms. In addition to (11), one further takes into account phenomenologically the pole term of the $\Lambda$ (1405) (abbreviated to $\Lambda^{\ast}$), which lies by about 30 MeV below the $K^{-}p$ threshold. The $\Lambda^{\ast}$ pole contribution is introduced by the Lagrangian for $K^{-}p\Lambda^{\ast}$ interaction:
@@
In Eq. (18), the dimensionless constants $d_{b}$ ($b=p,n,\Lambda,\Sigma^{-},\Xi^{-}$) of the range terms are related with $d_{p}\equiv(d_{1}+d_{2})m_{K}/2$ = $d_{\Xi^{-}}$, $d_{n}\equiv d_{1}m_{K}/2$ = $d_{\Sigma^{-}}$, $d_{\Lambda}=(d_{1}+5d_{2}/6)m_{K}/2$ in terms of $d_{1}$, $d_{2}$, and $m_{K}$. The mass difference between the $\Lambda^{\ast}$ and $p$ is denoted as $\delta M_{\Lambda^{\ast}p}\equiv M_{\Lambda^{\ast}}-M_{p}$, and $\gamma_{\Lambda^{\ast}}$ is introduced as the imaginary part of the inverse propagator for the $\Lambda^{\ast}$. The constants $d_{p}$, $d_{n}$, $g_{\Lambda^{\ast}}$, and $\gamma_{\Lambda^{\ast}}$ are determined to reproduce the $s$-wave on-shell $K$-$N$ scattering lengths including the pole contribution from $\Lambda^{\ast}$ martin1981 . 111 There has been a novel improvement on the $\bar{K}$-$N$ interaction by the SIDDHARTA experiment through the kaonic hydrogen $X$-rays siddharta2011 , while there is not a quantitatively large difference of the value of the $s$-wave $\bar{K}N$ scattering length between Ref. martin1981 and siddharta2011 . Therefore the conventional values in Ref. martin1981 are adopted. One obtains $g_{\Lambda^{\ast}}$ = 0.583, $\gamma_{\Lambda^{\ast}}$=12.4 MeV, $d_{p}$ = $0.351-\Sigma_{Kp}/m_{K}$, $d_{n}$ = $0.130-\Sigma_{Kn}/m_{K}$ (see Appendix X.2). As is the case with Ref.  fmmt1996 , the range terms and $\Lambda^{\ast}$-pole term are absorbed into the effective baryon mass as shown in Eq. (18).
It has become established by the use of chiral SU(3) dynamics that the $\Lambda^{\ast}$ is the $\bar{K}N$ bound state with the strong couplings such as $\bar{K}N\leftrightarrows\pi\Sigma$ ksw1995 ; or1998 ; lk2002 ; ihw2011 ; ro2000 . Properties of $\bar{K}$$N$ dynamics in the medium was studied in the chiral unitary approach by taking into account the self-consistent $\bar{K}N$ self-energy and the Pauli-blocking effects for the nucleon in the medium ro2000 . The two resonance pole structure of the $\Lambda^{\ast}$ has also been advocated om2001 ; joo2003 ; hj2012 . In this paper, one simply takes a conventional single pole picture for the $\Lambda^{\ast}$, since, as shown in Sec. VI.3, the $\Lambda^{\ast}$ contribution to the off-mass shell $K^{-}$ self-energy for $\rho_{\rm B}\gtrsim\rho_{0}$ is negligible, irrespective of details of the $\Lambda^{\ast}$ structure.